# Quasi-two-dimensional soliton in a self-repulsive spin-orbit-coupled dipolar binary condensate

### v1

@@ I Introduction
In mathematics and physics, a soliton is a nonlinear, one-dimensional (1D) self-reinforcing, localized wave packet that is strongly stable due to a balanced cancellation of attractive nonlinear and repulsive dispersive effects in the medium. A soliton can be found in diverse systems, for example, in a water wave, in a nonlinear medium nlm , in nonlinear optics x2 and in a Bose-Einstein condensate (BEC) x1 . Of these, a soliton in a BEC is especially interesting as it is a quantum system. In a quasi-1D set up, solitons have been created and studied in a BEC of 7Li r2a ; r2b and 85Rb r3 atoms, following the suggestion of a theoretical investigation r5 , using the mean-field Gross-Pitaevskii (GP) equation r6a ; r6b . There are numerous examples of similar quasi-1D solitons in BECs 38 ; 39 ; 40 ; 41 ; 42 ; 43 ; 44 ; 45 . Solitons have also been studied in binary BEC mixtures r7a ; r7b . However, stable solitons cannot be formed in two (2D) townes and three dimensions (3D) x1 ; x2 due to a collapse instability.
@@
Soon after the observation BEC of a BEC of 87Rb rb , 7Li li and 23Na na atoms in a laboratory, it was possible to observe and study a hyperfine spin-one multi-component spinor BEC of 23Na atoms exptspinor . Although it is not possible to have a natural spin-orbit (SO) coupling in electrically neutral BEC atoms, an artificial synthetic SO coupling can be created by tuning a pair of Raman lasers beams which couple the different hyperfine spin states thso ; exptsob ; sodip2 . Dresselhaus SOdre and Rashba SOras suggested two possible ways of creating such SO couplings. An equal mixture of Dresselhaus and Rashba SO couplings has been experimentally realized in a BEC of 87Rb exptso1 and 23Na na-solid atoms of $F_{z}=0,-1$ hyperspin components. The mean-field binary GP equation of such an SO-coupled BEC is coupled by the Pauli spin matrices and hence this BEC is called a pseudo spin-half BEC. Multi-component spinor BECs possess distinct properties and can have different excitations kita1 ; kita2 , which are not allowed in a single-component scalar BEC. For example, a quasi-1D quasi-1d , a quasi-2D quasi-2da ; quasi-2db or a 3D quasi-3d soliton can be stabilized in an SO-coupled binary BEC, which can have distinct spatial structure. Later, SO-coupled BECs were studied exptsob and an SO-coupled spin-one condensate of 87Rb atoms in hyperspin states $F_{z}=\pm 1,0$ was investigated exptsp1 . A spin-one SO-coupled multi-component BEC also sustains a quasi-1D quasi-1d1 or a quasi-2D quasi-2d1 or a 3D quasi-3d1 soliton, not possible in a single-component BEC townes .
@@
The studies of quasi-2D quasi-2d1 and 3D quasi-3d1 solitons in an SO-coupled nondipolar binary BEC indicate that the SO coupling introduces a stabilizing attractive interaction necessary for the formation of a self-bound soliton. All previous studies on SO-coupled nondipolar BEC solitons were performed on a self-attractive system with a negative atomic scattering length quasi-2da ; quasi-2db ; quasi-2d1 corresponding to an attraction among the atoms, which facilitates the formation of a soliton. A positive atomic scattering length corresponds to a repulsion among the atoms BEC . However, the scattering length in a BEC can often be varied continuosly from positive to negative, thus turning the system attractive from repulsive, by changing an external background magnetic field near a Feshbach resonance fesh ; fesh1 . In fact, this technique was used to change the scattering length of 85Rb atoms in the BEC from a positive value (repulsive interaction) to a negative value (attractive interaction) to facilitate the formation of a soliton. In this fashion a self-attractive BEC of 85Rb atoms was obtained and a quasi-1D soliton was created and studied r3 in this system in a laboratory. In a different front it was demonstrated malomed that it is possible to have a stable anisotropic soliton in a quasi-2D dipolar BEC with the polarization direction lying in the quasi-2D plane. There have also been many studies of quasi-1D solitons in dipolar BECs under different conditions 30 ; 31 ; 32 ; 33 .
@@
The long-range nonlocal anisotropic dipolar interaction sets up a different scenario malomed for the stabilization of a quasi-2D soliton, which will not be possible in the absence of a dipolar interaction townes . Many new phenomena are possible in a dipolar BEC, such as, angular collapse bohn , roton mode roton2 ; roton1 , quasi-2D asymmetric soliton malomed , quantum droplet drop , supersolidity in 1D ss1d and 2D ss2d , anisotropic expansion aaexp etc. The study of a dipolar BEC got new impetus after the experimental observation of dipolar BECs of 52Cr 1 ; 2 ; 6 , 166Er 7 , 168Er 8 , and 164Dy 9 ; 11 atoms, with large magnetic dipole moment. In contrast to previous studies kono ; adhisol ; quasi-2da ; quasi-2db of SO-coupled binary quasi-2D solitons in a self-attractive BEC, here we perform a comprehensive study of quasi-2D solitons in a binary Rashba or a Dresselhaus SO-coupled self-repulsive nondipolar and also dipolar BEC with positive scattering lengths, where the polarization direction of the dipolar atoms lies in the quasi-2D plane. The solitons in this system are bound due to a joint action of the SO coupling and the long-range dipolar interaction, in spite of the short-range contact repulsion in a self-repulsive binary BEC and new types of 2D solitons are expected in this system. We find drastic changes in soliton structure and density in an SO-coupled binary BEC soliton after the introduction of a weak dipolar interaction. There have been a few previous studies on different aspects of an SO-coupled dipolar BEC sodip1 ; sodip2 ; sodip3 ; sodip4 ; sodip5 ; sodip6 ; sodip7 .
@@
We find different types solitons in an SO-coupled quasi-2D binary nondipolar self-repulsive BEC. For a small value of SO-coupling strength parameter $\gamma$ ($0<\gamma\lessapprox 1$) there can be three types of degenerate solitons. The first type is a $(\mp 1,0)$-type multi-ring soliton, where the numbers in the parenthesis represent the angular momentum projections of the two components with the upper (lower) sign corresponding to Rashba (Dresselhaus) SO coupling. This state has 1/2 unit of angular momentum projection and is often called a half-quantum vortex state hpu . This half-quantum vortex state has been experimentally observed and studied in a topological superfluid 3He half-vortex . The second type is a stripe soliton with a spatially-periodic stripe pattern in density of the two components and no such stripe in the total density. Such stripe pattern in a pseudo spin-half SO-coupled BEC of 23Na atoms was experimentally observed na-solid . The third type is a circularly-asymmetric soliton. The component densities and total density of a circularly-asymmetric soliton are also circularly-asymmetric. For medium SO coupling ($\gamma\gtrapprox 1$), the aforementioned half-quantum vortex multi-ring soliton ceases to exist and we find three different types of degenerate solitons in a nondipolar binary BEC, e.g., a stripe soliton, a circularly-asymmetric soliton, and a square-lattice soliton with a spatially periodic 2D structure along both $x$ and $y$ directions.
@@
The introduction of a dipolar interaction in an SO-coupled BEC changes the scenario of soliton formation malomed . The dipolar interaction with the polarization direction in the quasi-2D $x$-$y$ plane breaks the circular symmetry in this plane. Consequently, the quasi-2D soliton gets stretched along the polarization direction. Because of this stretching, the square-lattice soliton ceases to exist and in this case we have only two types of degenerate solitons in a quasi-2D self-attractive or self-repulsive SO-coupled dipolar binary BEC: stripe solitons and circularly-asymmetric solitons. For all cases studied in this paper, the density and energy of the solitons are the same for Rashba and Dresselhaus SO couplings, although the two wave functions are different. In the presence of an SO coupling, magnetization is not conserved; However, the magnetization of most solitons studied in this paper is zero.
@@ II.1 Gross-Pitaevskii equation
The intraspecies ($V_{j},j=1,2$) and interspecies ($V_{12}$) interactions for two atoms at positions $\bf r$ and $\bf r^{\prime}$ are given by 1
@@
where $\bf R\equiv(r-r^{\prime})$ is the position vector joining the two atoms, $\mu_{0}$ is the permeability of free space, $\mu$ is the magnetic dipole moment of each atom, $\theta$ is the angle made by the vector $\bf R$} with the polarization $y$ direction, $a_{j}$ is the intraspecies scattering length, and $a_{12}$ is the interspecies scattering length. To compare the dipolar and contact interactions, the dipolar interaction is expressed in terms of the dipolar length, defined by 1
@@
This dipolar binary BEC of two components $j=1,2$ is described by the following set of equations 1 ; bao
@@
where $\partial_{t}\equiv\partial/\partial t$, $N_{j}$ and $N_{k}$ are the number of atoms in the two components ($N=N_{1}+N_{2}$). The dipolar atoms are polarized along the $x$ direction. Equations (II.1) can be cast in the following dimensionless form if we scale lengths in units of $l_{0}=\sqrt{\hbar/m\omega_{z}}$, time units of $t_{0}=1/\omega_{z}$, density in units of $l_{0}^{-2}$, and energy in units of $\hbar\omega_{z}$
@@
We consider a quasi-2D binary dipolar BEC with a strong trap along the $z$ direction. We assume that the dynamics of the BEC in the $z$ direction is confined in the ground state
@@
The component wave function of the BEC can be written as
@@
where $\psi_{i}({\boldsymbol{\rho}},t)$ is the wave function of the quasi-2D BEC confined in the $x$-$y$ plane. The relevant dynamics in the $x$-$y$ plane, obtained by integrating out the $z$ coordinate 2d-3d , leads to the following set of GP equations for the two components bec2015 ; bao
@@
The time-independent version of Eqs. (II.1), appropriate for stationary solutions, can be derived from the energy functional
@@
$+2\gamma\{\psi_{1}^{*}(\partial_{x}-{\mbox{i}}\eta\partial_{y})\psi_{2}-\psi_{2}^{*}(\partial_{x}+{\mbox{i}}\eta\partial_{y})\psi_{1}\}\Big]\,.$ (17)
@@ II.2 Analytic Consideration
with two equivalent ground states. For the first state, $m=0$ and the second component has a phase $\exp({\mbox{i}}\eta\theta)$ relative to the first component. This state is of type $(0,\eta)$ or $(0,\pm 1)$ for Rashba (upper sign) or Dresselhaus (lower sign) SO coupling. The second state has $m=-\eta$ with the first component carrying a phase $\exp(-{\mbox{i}}\eta\theta)$ relative to the second component. This state is of type $(-\eta,0)$ or $(\mp 1,0)$ for Rashba or Dresselhaus SO coupling. Of these two types of equivalent states, in this study we consider only the state of type $(\mp 1,0)$. These states have angular momentum projection $m=0$ and spin $s=1/2$, hence the total angular momentum projection $j_{z}=m+s_{z}=\pm 1/2$ and these states are often called half-quantum vortex states. However, if we consider larger values of $|m|>1$ in Eq. (20), we can have excited states of type $(\mp 2,\mp 1),(\mp 3,\mp 2)$ etc. of larger energy. In an actual nondipolar SO-coupled uniform BEC for small $\gamma$ the infinite ground state of type $(\mp 1,0)$ becomes a localized multi-ring soliton.
@@ III Numerical Result
To solve the binary equations (II.1) numerically, we propagate these in time by the split-time-step Crank-Nicolson discretization scheme bec2009 . As we require to solve the GP equation in presence of both SO-coupling and dipolar interaction, we needed to combine the Open Multiprocessing Programs for solving the dipolar bec2015 ; bec2023 and SO-coupled bec2021 GP equations. We employ space steps of $dx=dy=0.1$ and a time step of $dt=dx\times dy\times 0.1$ for imaginary-time propagation. In all calculations of stationary states, we employ imaginary-time approach with the conservation of total normalization
@@
is not a good quantum number in the presence of spin-mixing SO coupling and is left to freely evolve during time propagation to attain a final converged value consistent with the parameters of the problem. In the presence of SO coupling the number of atoms in each component is not conserved. The converged value of the magnetization $M$ was found to be negligibly small in most cases for values of $\gamma$ considered in this paper indicating an almost equal number of atoms in the two components. For $\gamma=0$, the magnetization is strictly equal to zero.
@@
In an actual experiment, by carefully ramping up the intensity of the Raman beams used to create the SO coupling and adjusting the detuning, the system undergoes a phase transition into the stripe phase. The competition between the interaction energy and the SOC term in the Hamiltonian leads to a ground state where the atoms condense in a superposition of two finite-momentum states, spontaneously breaking the translational symmetry of space. The multi-ring phase can be achieved by imprinting an appropriate vortex by a laser on the BEC.

### v2

@@ I Introduction
In mathematics and physics, a soliton is a nonlinear, one-dimensional (1D) self-reinforcing, localized wave packet that is strongly stable due to a balanced cancellation of attractive nonlinear and repulsive dispersive effects in the medium. A soliton can be found in diverse systems, for example, in a water wave, in a nonlinear medium nlm , in nonlinear optics x2 and in a Bose-Einstein condensate (BEC) x1 . Of these, a soliton in a BEC is especially interesting as it is a quantum system. In a quasi-one-dimensional (quasi-1D) set up, solitons have been created and studied in a BEC of 7Li r2a ; r2b and 85Rb r3 atoms, following the suggestion of a theoretical investigation r5 , using the mean-field Gross-Pitaevskii (GP) equation r6a ; r6b . There are numerous examples of similar quasi-1D solitons in BECs 38 ; 39 ; 40 ; 41 ; 42 ; 43 ; 44 ; 45 . Solitons have also been studied in binary BEC mixtures r7a ; r7b . However, stable solitons cannot be formed in two (2D) townes and three dimensions (3D) x1 ; x2 due to a collapse instability.
@@
Soon after the observation BEC of a BEC of 87Rb rb , 7Li li and 23Na na atoms in a laboratory, it was possible to observe and study a hyperfine spin-one multi-component spinor BEC of 23Na atoms exptspinor . Although it is not possible to have a natural spin-orbit (SO) coupling in electrically neutral BEC atoms, an artificial synthetic SO coupling can be created in neutral atoms by tuning a few Raman laser beams, which couple the different hyperfine spin states thso ; exptsob ; sodip2 . Dresselhaus SOdre and Rashba SOras suggested two possible ways of creating such SO couplings. An equal mixture of Dresselhaus and Rashba SO couplings has been experimentally realized in a BEC of 87Rb exptso1 and 23Na na-solid atoms of $F_{z}=0,-1$ hyperspin components. The mean-field binary GP equation of such an SO-coupled BEC is coupled by the Pauli spin matrices and hence this BEC is called a pseudo spin-half BEC. Multi-component spinor BECs possess distinct properties and can have different excitation kita1 ; kita2 , which are not allowed in a single-component scalar BEC. For example, a quasi-1D quasi-1d , a quasi-two-dimensional (quasi-2D) quasi-2da ; quasi-2db or a 3D quasi-3d soliton can be stabilized in an SO-coupled binary BEC, which can have distinct spatial structure. Later, SO-coupled BECs were studied exptsob and an SO-coupled spin-one condensate of 87Rb atoms in hyperspin states $F_{z}=\pm 1,0$ was investigated exptsp1 . A spin-one SO-coupled multi-component BEC also sustains a quasi-1D quasi-1d1 or a quasi-2D quasi-2d1 or a 3D quasi-3d1 soliton, not possible in a single-component BEC townes .
@@
The studies of quasi-2D quasi-2d1 and 3D quasi-3d1 solitons in an SO-coupled nondipolar binary BEC indicate that the SO coupling introduces a stabilizing attractive interaction necessary for the formation of a self-bound soliton. All previous studies on SO-coupled nondipolar BEC solitons were performed on a self-attractive system with a negative atomic scattering length quasi-2da ; quasi-2db ; quasi-2d1 corresponding to an attraction among the atoms, which facilitates the formation of a soliton. A positive atomic scattering length corresponds to a repulsion among the atoms BEC . However, the scattering length in a BEC can often be varied continuously from positive to negative, thus making the system attractive from repulsive, by changing an external background magnetic field near a Feshbach resonance fesh ; fesh1 . In fact, this technique was used to change the scattering length of 85Rb atoms in the BEC from a positive value to a negative value to obtain a self-attractive BEC of 85Rb atoms and a quasi-1D soliton was created and studied r3 in this system in a laboratory. In a different front it was demonstrated malomed that it is possible to have a stable anisotropic soliton in a quasi-2D dipolar BEC with the polarization direction lying in the quasi-2D plane. There have also been many studies of quasi-1D solitons in dipolar BECs under different conditions 30 ; 31 ; 32 ; 33 .
@@
The long-range nonlocal anisotropic dipolar interaction sets up a different scenario malomed for the stabilization of a quasi-2D soliton, which will not be possible in the absence of a dipolar interaction townes . Many new phenomena are possible in a dipolar BEC, such as, angular collapse bohn , roton mode roton2 ; roton1 , quasi-2D asymmetric soliton malomed , quantum droplet drop , supersolidity in 1D ss1d and 2D ss2d , anisotropic expansion aaexp etc. The study of a dipolar BEC got new impetus after the experimental observation of dipolar BECs of 52Cr 1 ; 2 ; 6 , 166Er 7 , 168Er 8 , and 164Dy 9 ; 11 atoms, with large magnetic dipole moment.
In contrast to previous studies kono ; adhisol ; quasi-2da ; quasi-2db of SO-coupled binary quasi-2D solitons in a self-attractive BEC, here we perform a comprehensive study of quasi-2D solitons in a binary Rashba or a Dresselhaus SO-coupled self-repulsive nondipolar and also dipolar BEC with positive scattering lengths, where the polarization direction of the dipolar atoms lies in the quasi-2D plane. The solitons in this system are bound due to a joint action of the SO coupling and the long-range dipolar interaction, in spite of the short-range contact repulsion in a self-repulsive binary BEC and new types of 2D solitons are expected in this system. We find drastic changes in soliton structure and density in an SO-coupled binary BEC soliton after the introduction of a weak dipolar interaction. There have been a few previous studies on different aspects of an SO-coupled dipolar BEC sodip1 ; sodip2 ; sodip3 ; sodip4 ; sodip5 ; sodip6 ; sodip7 ; yuka ; yuka1 .
@@
We find different types solitons in an SO-coupled quasi-2D binary nondipolar self-repulsive BEC. For a small value of SO-coupling strength parameter $\gamma$ ($0<\gamma\lessapprox 1$) there can be three types of degenerate solitons. The first type is a $(\mp 1,0)$-type multi-ring soliton, where the numbers in the parenthesis represent the angular momentum projections of the two components with the upper (lower) sign corresponding to Rashba (Dresselhaus) SO coupling. This state has 1/2 unit of angular momentum projection and is often called a half-quantum vortex state hpu . This half-quantum vortex state has been experimentally observed and studied in a topological superfluid 3He half-vortex ; hv . The second type is a stripe soliton with a spatially-periodic stripe pattern in density of the two components and no such stripe in the total density. Such stripe pattern in a pseudo spin-half SO-coupled BEC of 23Na atoms was experimentally observed na-solid . The third type is a circularly-asymmetric soliton. The component densities and total density of a circularly-asymmetric soliton are also circularly-asymmetric. For medium SO coupling ($\gamma\gtrapprox 1$), the aforementioned half-quantum vortex multi-ring soliton ceases to exist and we find three different types of degenerate solitons in a nondipolar binary BEC, e.g., a stripe soliton, a circularly-asymmetric soliton, and a square-lattice soliton with a spatially periodic 2D structure along both $x$ and $y$ directions.
@@
The introduction of a dipolar interaction in an SO-coupled BEC changes the scenario of soliton formation malomed . The dipolar interaction, with the polarization direction in the quasi-2D $x$-$y$ plane, breaks the circular symmetry in this plane. Consequently, the quasi-2D soliton gets stretched along the polarization direction. Because of this stretching, the square-lattice soliton ceases to exist and in this case we have only two types of degenerate solitons in a quasi-2D self-attractive or self-repulsive SO-coupled dipolar binary BEC: stripe solitons and circularly-asymmetric solitons. For all cases studied in this paper, the density and energy of the solitons are the same for Rashba and Dresselhaus SO couplings, although the two wave functions are different. In the presence of an SO coupling, magnetization is not conserved. However, the magnetization of most solitons studied in this paper is zero.
@@ II.1 Gross-Pitaevskii equation
The interaction between the two atoms has the following two parts: the contact interaction $\delta({\bf R})$ and the nonlocal asymmetric dipolar interaction $V_{\mathrm{dd}}({\bf R}).$ The intraspecies ($V_{j},j=1,2$) and interspecies ($V_{12}$) interactions for two atoms at positions $\bf r$ and $\bf r^{\prime}$ are given by 1
@@
where $\bf R\equiv(r-r^{\prime})$ is the position vector joining the two atoms, $\mu_{0}$ is the permeability of free space, $\mu$ is the magnetic dipole moment of each atom, $\theta$ is the angle made by the vector $\bf R$} with the polarization $y$ direction, $a_{j}$ is the intraspecies scattering length, and $a_{12}$ is the interspecies scattering length. The scattering length measures the strength of contact interaction in an ultra-dilute BEC. To compare the dipolar and contact interactions, the dipolar interaction is expressed in terms of the dipolar length, defined by 1
@@
The dipolar binary BEC of two components $j=1,2$ is described by the following set of equations 1 ; bao
@@
where $\partial_{t}\equiv\partial/\partial t$, $N_{j}$ and $N_{k}$ are the number of atoms in the two components ($N=N_{1}+N_{2}$). The dipolar atoms are polarized along the $y$ axis. Equations (II.1) can be cast in the following dimensionless form if we scale lengths in units of $l_{0}=\sqrt{\hbar/m\omega_{z}}$, time units of $t_{0}=1/\omega_{z}$, density in units of $l_{0}^{-2}$, and energy in units of $\hbar\omega_{z}$
@@
We consider a quasi-2D binary dipolar BEC with a strong trap along the $z$ direction. We assume that the dynamics of the BEC in the $z$ direction is frozen in the ground state
@@
and the relevant dynamics of the dipolar BEC will be confined in the $x$-$y$ plane. In this case it is possible to integrate out the $z$ variable and write a set of binary coupled equations for the relevant dynamics in the $x$-$y$ plane. The component wave function of the BEC can be written as
@@
where $\psi_{i}({\boldsymbol{\rho}},t)$ is the wave function of the quasi-2D BEC confined in the $x$-$y$ plane. To obtain the relevant dynamics, we substitute Eq. (10) into Eq. (8), multiply the resultant equation by $\psi_{B}(z)$ and integrate over $z$ 2d-3d ; bec2015 . To avoid the difficulty associated with the integration over the divergent dipolar interaction in the configuration space, that integral was evaluated in the momentum space. This procedure 2d-3d leads to the following set of GP equations for the two components bec2015 ; bao
@@
Equations (II.1) can be derived from the energy functional
@@
$+2\gamma\{\psi_{1}^{*}(\partial_{x}-{\mbox{i}}\eta\partial_{y})\psi_{2}-\psi_{2}^{*}(\partial_{x}+{\mbox{i}}\eta\partial_{y})\psi_{1}\}\Big]\,$, (17)
using the variational rule BEC
$\mathrm{i}\partial_{t}\psi_{j}=\frac{\delta E[\psi_{j}]}{\delta\psi_{j}}$ (18)
@@ II.2 Analytic Consideration
with two equivalent ground states. For the first state, $m=0$ and the second component has a phase $\exp({\mbox{i}}\eta\theta)$ relative to the first component. This state is of type $(0,\eta)$ or $(0,\pm 1)$ for Rashba (upper sign) or Dresselhaus (lower sign) SO coupling. The second state has $m=-\eta$ with the first component carrying a phase $\exp(-{\mbox{i}}\eta\theta)$ relative to the second component. This state is of type $(-\eta,0)$ or $(\mp 1,0)$ for Rashba or Dresselhaus SO coupling. Of these two types of equivalent states, in this study we consider only the state of type $(\mp 1,0)$. These states have angular momentum projection $m=0$ and spin $s=1/2$, hence the total angular momentum projection $j_{z}=m+s_{z}=\pm 1/2$ and these states are often called half-quantum vortex states hpu ; hv . However, if we consider larger values of $|m|>1$ in Eq. (21), we can have excited states of type $(\mp 2,\mp 1),(\mp 3,\mp 2)$ etc. of larger energy. In an actual nondipolar SO-coupled uniform BEC for small $\gamma$ the infinite ground state of type $(\mp 1,0)$ becomes a localized multi-ring soliton.
@@ III Numerical Result
To solve the binary equations (II.1) numerically, we propagate these in time by the split-time-step Crank-Nicolson discretization scheme bec2009 . As we require to solve the GP equation in the presence of both SO-coupling and dipolar interaction, we needed to combine the Open Multiprocessing Programs for solving the dipolar bec2015 ; bec2023 and SO-coupled bec2021 GP equations. We employ space steps of $dx=dy=0.1$ and a time step of $dt=dx\times dy\times 0.1$ for imaginary-time propagation. In all calculations of stationary states, we employ imaginary-time approach with the conservation of total normalization
@@
is not a good quantum number in the presence of spin-mixing SO coupling and is left to freely evolve during time propagation to attain a final converged value consistent with the parameters of the problem. In the presence of SO coupling the number of atoms in each component is not conserved due to the possibility of a spontaneous transfer of atoms from one spin state to the other. The converged value of the magnetization $M$ was found to be negligibly small in most cases for values of $\gamma$ considered in this paper indicating an almost equal number of atoms in the two components. For $\gamma=0$, the magnetization is strictly equal to zero provided that the initial magnetization is zero.
@@
In an actual experiment 14 , by carefully ramping up the intensity of the Raman beams used to create the SO coupling and adjusting the detuning, the system undergoes a phase transition into the stripe phase. The competition between the interaction energy and the SO-coupling term in the Hamiltonian leads to a ground state where the atoms condense in a superposition of two finite-momentum states, spontaneously breaking the translational symmetry of space. The multi-ring phase can be achieved by imprinting an appropriate vortex by a laser on the BEC.